OpenStudy (anonymous):

Which expression is the partial fraction decomposition of  3x+7/ (x+2)^2

OpenStudy (anonymous):

Can someone please help me ? :(

OpenStudy (anonymous):

sure you need to know it is going to look like \[\frac{A}{x+2}+\frac{B}{(x+2)^2}\] first right?

OpenStudy (anonymous):

is that part clear? if so then we can do it quickly

OpenStudy (anonymous):

So would it be 3 over x+7 and x over x+2 ?

OpenStudy (anonymous):

no we need to find \(A\) and \(B\)

OpenStudy (anonymous):

your denominator is \[(x+2)^2\] right?

OpenStudy (anonymous):

So it would just be A over x+7 ? and then B over x+2 ^2 ?

OpenStudy (anonymous):

there is no \(x+7\) in it as i said, it must look like \[\frac{A}{x+2}+\frac{B}{(x+2)^2}\]for some number A and B

OpenStudy (jdoe0001):

hmmm do you know what "partial decomposition" is?

OpenStudy (anonymous):

No. this online class im taking is honestly not teaching me anything

OpenStudy (jdoe0001):

partial = split in two or more parts fraction = well, an expression with numerator and denominator decomposition = to decompose, or expand over

OpenStudy (jdoe0001):

so.. what satellite73 is pointing out is that, you have a SUM RESULT, and you're asked to produce 2 fractions that give that result

OpenStudy (anonymous):

So the equation should look like the picture he posted. okay so after that how would i solve it ?

OpenStudy (jdoe0001):

http://www.purplemath.com/modules/partfrac.htm <---- this should cover it I'd think

OpenStudy (anonymous):

Im sorry but this isnt helping me . Im not understanding the way there explaining it

OpenStudy (anonymous):

I got it : ) . I got the answer right thank you. Do you mind helping me on another question ?

OpenStudy (jdoe0001):

sure.. just post anew... thus if I dunno, someone else may